OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 3x - 4y = -12.

OpenStudy (anonymous):

A. intercept = (3, 0); y-intercept = (0, -4) B. x-intercept = (-4, 0); y-intercept = (0, 3) C. intercept = (0, 3); y-intercept = (-4, 0) D. intercept = (0, -4); y-intercept = (3, 0)

OpenStudy (anonymous):

x=0 -4y=-12 -12/-4 y=o,3 y=0 3x=-12 -12/-3 x=-4,0

OpenStudy (anonymous):

b